OpenStudy (anonymous):

Find the capacitance of a capacitor that stores 370μJ when the potential difference across its plates is 150V .

OpenStudy (souvik):

Energy=1/2CV^2

OpenStudy (souvik):

here C=capacitance and V=voltage difference

OpenStudy (souvik):

370*10^-6 =1/2*C*150^2